### v3

@@ Abstract
In this work, we provide data stream algorithms that compute optimal splits in decision tree learning. In particular, given a data stream of observations $x_{i}$ and their labels $y_{i}$, the goal is to find the optimal split $j$ that divides the data into two sets such that the mean squared error (for regression) or misclassification rate and Gini impurity (for classification) is minimized. We provide several fast streaming algorithms that use sublinear space and a small number of passes for these problems. These algorithms can also be extended to the massively parallel computation model. Our work, while not directly comparable, complements the seminal work of Domingos-Hulten (KDD 2000) and Hulten-Spencer-Domingos (KDD 2001).
@@ 1 Introduction
A key step in building a decision tree is to find optimal splits. Denote $[N]$ as the set $\{1,2,\ldots,N\}$. The data consists of observations $x_{1},x_{2},\ldots,x_{m}\in[N]$ and their real-valued labels $y_{1},y_{2},\ldots,y_{m}\in[0,M]$ (for regression) or binary labels $y_{1},y_{2},\ldots,y_{m}\in{-1,+1}$ (for classification). Informally, we want to find the optimal split $j\in[N]$ to partition the data into two sets, $\{x_{i}\leq j\}$ and $\{x_{i}>j\}$, such that the mean squared error (MSE) for regression or the misclassification rate for classification is minimized.
@@
In the case of regression, we assume that $y_{i}$ is in the range $[0,M]$ and each $y_{i}$ can be stored in one machine word of size $O(\log N)$ bits.
In many applications, the data is too large to fit in the main memory of a single machine. One often streams the data from disk or network and processes the data in a streaming manner in sublinear space and ideally in a small number of passes. In some applications, a small number of passes is allowed.
@@
Our second main result is as follows.
@@ C.2 $\tilde{O}\left(1/\epsilon^{2}\right)$ $1$ $L_{Gini}(j)\leq\operatorname{OPT}+\epsilon$
Previously, a very popular work by Domingos and Hulten [7]  [^4] Their paper was cited more than 3000 times according to Google Scholar. proposed a decision tree learning algorithm for data streams. In their study, they considered the scenario where one has an infinite stream of observations where each observation is independently and identically distributed (i.i.d.). They quantified the number of samples required to find the optimal split and provided a bound on the probability that a sample will follow a different path in the decision tree than it would in a tree built using the entire data set under some assumptions. It is worth noting that the work of Domingos and Hulten [7] did not consider the regression problem. Subsequently, Hulten, Spencer, and Domingos [18] provided a heuristic to build a decision tree based on most recent data in the stream. Some variant as well as improvements for this model are discussed in [18, 20, 27].
Compared to the work in [7, 18, 20, 27], our work does not assume that the data stream is infinite. We also do not make the i.i.d. assumption. This is useful in scenarios such as when the data stream is collected in a time-dependent manner and the distribution of the data changes over time. Furthermore, we also consider the regression problem.
Essentially, our work builds a depth-1 tree for both classification and regression in the streaming and MPC models. Thus, to construct a deeper tree, we can repeatedly apply our algorithm to build a deeper tree using any decision tree learning algorithm of interest such as ID3 [26], C4.5 [26], or CART [2] with additional passes through the data stream. We remark that the problem of computing the optimal decision tree is generally NP-Hard [19, 24].
If the observations and labels are i.i.d., our algorithm can be extended to build a tree by utilizing batches of observations to determine the optimal splits as the tree grows. In this case, our approach also accommodates the regression problem. Moreover, our work quantifies the “error” at each node in the tree, offering an advantage when one does not wish to benchmark against a specific tree learning algorithm as done in [7, 18, 20, 27].
@@ 1 When $x_{i},y_{i}$ arrive, update the following quantities:
We need to compute $\sum_{t\leq j}A_{t}$, $\sum_{t>j}A_{t},\sum_{t\leq j}B_{t}$, $\sum_{t>j}B_{t}$, $\sum_{t\leq j}C_{t}$, and $\sum_{t>j}C_{t}$ for all $t$ in $O(D)$ time. This can be easily done using dynamic programming (e.g., $\sum_{t\leq j}B_{t}=(\sum_{t<j}B_{t})+B_{j}$ and $\sum_{t>j}B_{t}=(\sum_{t>j+1}B_{t})+B_{j+1}$). This allows us to compute $\mu(j)$, $\gamma(j)$, and $L(j)$ for all distinct values $j$ of $\{x_{1},x_{2},\ldots,x_{m}\}$.
@@
$=\frac{1}{m}\left(\sum_{t\leq j}C_{t}+\mu(j)^{2}A_{t}-2\mu(j)B_{t}+\sum_{t>j}C_{t}+\gamma(j)^{2}A_{t}-2\gamma(j)B_{t}\right)$
@@
The idea is to find a set of $\approx 1/\epsilon$ candidate split points $S$ such that the number of observations between any two consecutive split points is at most $\epsilon m$. Then, we can compute $L(j)$ for all $j\in S$ and return the split point $j$ with the smallest $L(j)$. This will gives a $j$ such that $L(j)\leq\operatorname{OPT}+O(\epsilon)$.
@@ 6 For each $t$ such that $j_{t},j_{t+1}\in S$, using dynamic programming, compute
We next establish the approximation guarantee. Note that if we sample some $x_{i}=j$, then we add both $j$ and $j-1$ to the candidate set. Suppose the optimal split $j^{\star}\in S$ then we are done since $L(j^{\star})$ will be computed in the second pass. Otherwise, there exists $t$ such that $j_{t}<j^{\star}<j_{t+1}$. This implies that $j_{t}<j_{t+1}-1$; hence, it must be the case that $j_{t+1}$ was added to $S$ because we sample some $x_{i}=j_{t+2}$ and we did not sample any $x_{i}=j_{t+1}$. Therefore, the algorithm did not sample any observation in $(j_{t},j_{t+1}]\supset(j^{\star},j_{t+1}]$. This implies
@@ 7             Make another pass over the stream and compute
if $\texttt{Error}_{left}(j)\leq z_{left}$ and $\text{Error}_{right}(j)\leq z_{right}$ then
@@ 5       Define the estimate
7       if there are fewer than $T=K/\epsilon^{2}\cdot\log N$ samples in $S_{2_{j}}$ then

### v4

@@ Abstract
In this work, we present data stream algorithms to compute optimal splits for decision tree learning. In particular, given a data stream of observations $x_{i}$ and their corresponding labels $y_{i}$, without the i.i.d. assumption, the objective is to identify the optimal split $j$ that partitions the data into two sets, minimizing the mean squared error (for regression) or the misclassification rate and Gini impurity (for classification). We propose several efficient streaming algorithms that require sublinear space and use a small number of passes to solve these problems. These algorithms can also be extended to the MapReduce model. Our results, while not directly comparable, complements the seminal work of Domingos-Hulten (KDD 2000) and Hulten-Spencer-Domingos (KDD 2001).
@@ 1 Introduction
A key step in building a decision tree is to find optimal splits. Denote $[N]$ as the set $\{1,2,\ldots,N\}$. The data consists of observations $x_{1},x_{2},\ldots,x_{m}\in[N]$ and their real-valued labels $y_{1},y_{2},\ldots,y_{m}\in[0,M]$ (for regression) or binary labels $y_{1},y_{2},\ldots,y_{m}\in\{-1,+1\}$ (for classification). Informally, we want to find the optimal split $j\in[N]$ to partition the data into two sets, $\{x_{i}\leq j\}$ and $\{x_{i}>j\}$, such that the mean squared error (MSE) for regression or the misclassification rate for classification is minimized.
@@
In the case of regression, we assume that $y_{i}\in[0,M]$ and each $y_{i}$ has finite precision that can be stored in one machine word of size $O(\log N)$ bits.
In many applications, the data is too large to fit into the main memory of a single machine. Often, the data is streamed from disk or over a network while using sublinear space, ideally in a single pass. In some cases, a small number of passes is allowed.
@@
Our results for classification are as follows.
@@ C.2 $\tilde{O}\left(1/\epsilon^{2}\right)$ $1$ $L_{Gini}(j)\leq\operatorname{OPT}+\epsilon$
Previously, a very popular work by Domingos and Hulten [11]  [^4] Their paper was cited more than 3000 times according to Google Scholar. proposed a decision tree learning algorithm for data streams. In their study, they considered the scenario where one has an infinite stream of observations where each observation is independently and identically distributed (i.i.d.). They quantified the number of samples required to find the optimal split and provided a bound on the probability that a sample will follow a different path in the decision tree than it would in a tree built using the entire data set under some assumptions. It is worth noting that the work of Domingos and Hulten [11] did not consider the regression problem.
Subsequently, various heuristics, implementations, and applications of streaming decision trees have been developed. Providing a comprehensive list of related work is challenging. Notable examples include Gomes et al. [17], who introduced an adaptive random forest algorithm for classification in evolving data streams. Additionally, C++ and Python systems [3, 29] have been developed to implement decision trees for data streams, which have seen extensive commercial use. Another popular application of streaming decision trees is mining concept drift [34, 6]. Other notable works include [23, 24, 2, 32, 4, 27].
Compared to these works, our approach does not assume an infinite data stream. Additionally, we do not rely on the i.i.d. assumption, making our method more applicable to scenarios where data streams are collected in a time-dependent manner and the data distribution evolves over time. Furthermore, we also consider the regression problem. Our algorithms come with provable guarantees.
@@ 1 When $x_{i},y_{i}$ arrive, update the following quantities:
We need to compute $\sum_{t\leq j}A_{t}$, $\sum_{t>j}A_{t},\sum_{t\leq j}B_{t}$, $\sum_{t>j}B_{t}$, $\sum_{t\leq j}C_{t}$, and $\sum_{t>j}C_{t}$ for all $t$ in $O(D)$ time. This can be easily done using prefix-sum dynamic programming (e.g., $\sum_{t\leq j}B_{t}=(\sum_{t<j}B_{t})+B_{j}$ and $\sum_{t>j}B_{t}=(\sum_{t>j+1}B_{t})+B_{j+1}$). This allows us to compute $\mu(j)$, $\gamma(j)$, and $L(j)$ for all distinct values $j$ of $\{x_{1},x_{2},\ldots,x_{m}\}$.
@@
$=\frac{1}{m}\left(\sum_{t\leq j}(C_{t}+\mu(j)^{2}A_{t}-2\mu(j)B_{t})+\sum_{t>j}(C_{t}+\gamma(j)^{2}A_{t}-2\gamma(j)B_{t})\right)$
@@
The idea is to find a set of $\approx 1/\epsilon$ candidate split points $S$ such that the number of observations between any two consecutive split points is at most $\epsilon m$. Then, we can compute $L(j)$ for all $j\in S$ and return the split point $j$ with the smallest $L(j)$. This gives a $j$ such that $L(j)\leq\operatorname{OPT}+O(\epsilon)$.
@@ 6 For each $t$ such that $j_{t},j_{t+1}\in S$, using dynamic programming, compute
We next establish the approximation guarantee. Note that if we sample some $x_{i}=j$, then we add both $j$ and $j-1$ to the candidate set. Suppose the optimal split $j^{\star}\in S$ then we are done since $L(j^{\star})$ will be computed in the second pass. Otherwise, there exists $t$ such that $j_{t}<j^{\star}<j_{t+1}$. This implies that $j_{t}<j_{t+1}-1$; hence, it must be the case that $j_{t+1}$ was added to $S$ because we sample some $x_{i}=j_{t+2}$ and we did not sample any $x_{i}=j_{t+1}$. Therefore, the algorithm did not sample any observation in $(j_{t},j_{t+1}]\supset(j^{\star},j_{t+1}]$. With $a=j_{\star}+1$ and $b=j_{t+1}$, this implies
@@ 7             Make another pass over the stream and compute
if $\texttt{Error}_{left}(j)\leq z_{left}$ and $\texttt{Error}_{right}(j)\leq z_{right}$ then
@@ 5       Define the estimate
7       if there are fewer than $T=K/\epsilon^{2}\cdot\log N$ samples in $S_{2,j}$ then